OpenStudy (anonymous):

multiply 5(sqrt)3(2(sqrt)3 + 6 (sqrt)19)

OpenStudy (anonymous):

How did you go with this one?

OpenStudy (anonymous):

okay first i distribute the 5 into 2 and 6 cause i remember outside with the out and in side with the in so now im stuck

OpenStudy (anonymous):

okay good so you have \[10\sqrt{3}\sqrt{3} + 30\sqrt{3}\sqrt{19}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so its 30(sqrt)3

OpenStudy (anonymous):

+ 90(sqrt)19?

OpenStudy (anonymous):

sorry with the first term we had 10 sqrt(3)*sqrt(3) remember that sqrt(a)*sqrt(a)=a so what will your first term be?

OpenStudy (anonymous):

10(sqrt)3

OpenStudy (anonymous):

it will be 10*3=30 do you see this?

OpenStudy (anonymous):

30(sqrt)3

OpenStudy (anonymous):

no it will just be 30 \[10*\sqrt{3}*\sqrt{3}=10*3=30\]

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

so with the 2nd term

OpenStudy (anonymous):

30 + 30(sqrt)57?

OpenStudy (anonymous):

correct!!!

OpenStudy (anonymous):

thank you!